# Collaborative Object Handover in a Robot Crafting Assistant

### v1

@@ Abstract
Robots are increasingly working alongside people, delivering food to patrons in restaurants or helping workers on assembly lines. These scenarios often involve object handovers between the person and the robot. To achieve safe and efficient human-robot collaboration (HRC), it is important to incorporate human context in a robot’s handover strategies. Therefore, in this work, we develop a collaborative handover model trained on human teleoperation data collected in a naturalistic crafting task. To evaluate the performance of this model, we conduct cross-validation experiments on the training dataset as well as a user study in the same HRC crafting task. The handover episodes and user perceptions of the autonomous handover policy were compared with those of the human teleoperated handovers. While the cross-validation experiment and user study indicate that the autonomous policy successfully achieved collaborative handovers, the comparison with human teleoperation revealed avenues for further improvements.
I Introduction
@@
Object handover is a common task in human-robot collaboration (HRC), in which an item is transferred between a human and a robot [1]. To develop an autonomous human-robot handover policy, existing approaches either react to pre-defined triggers (e.g., specific hand gestures [2]), or train machine learning models using data collected from isolated handover episodes between two people or in human-robot pairs (e.g., [3, 4]). However, using pre-defined triggers imposes learning costs for the person and interrupts the flow of activities, requiring additional physical and mental effort for the person to explicitly command the robot [5]. Further, as handovers are context-dependent [1], training from isolated episodes limits the resulting model’s applicability in handovers that occur in naturalistic contexts with pre- and post-handover tasks.
@@
There has been growing interest in the social-interactive aspects of handovers in addition to the control and trajectory planning aspects, focusing on how a person and a robot may communicate and coordinate during handovers [6, 7]. While these studies investigated communicating information such as intention or hesitation during individual episodes of object transfer, how people and robots collaborate in a complex task involving multiple handovers remains largely under explored. Handovers in a broader task context require providing objects at the right place and the right time, when the person may be engaged in the task and not giving explicit cues to the robot. Further, research on human behaviours in handovers largely focuses on hand locations, while additional informative behaviours such as gaze or whole body movements are under explored [8]. Evaluation of human-robot handover policies is often conducted in a short-term, handover-centred design without incorporating human and task contexts, e.g., focusing on the smoothness of the robot’s motions and trajectories [9]. This results in a gap between a handover policy’s outcomes in isolated episodes compared to when multiple handovers are serving an overall task goal. For example, a policy that emphasises optimising the instantaneous ergonomics risk factors in individual handovers may become over-assistive and reduce worker well-being in repeated interactions [10].
Top half of the image shows a photo of the experimental layout, with the work space where participants sit on the left and storage space where the experimenter sits on the right. The Fetch robot is moving between these two desks. A partially built birdhouse can be seen on top of the work desk. A tripod with a camera on top stands next to the work desk. A storage box can be seen next to the storage space. The robot also has a camera on its head. Below the photo, the state transitions during a handover episode are shown, with three timing adaptations and two handover location adaptations highlighted at key state transitions.
@@
Therefore, we are motivated to understand human-robot handovers in a naturalistic task context where they serve a collaborative goal. In particular, we develop an autonomous handover policy, as shown in Figure 1, trained with the Functional And Creative Tasks Human-Robot Collaboration dataset (the FACT HRC dataset) [11]. In the FACT HRC dataset, a mobile manipulator robot was teleoperated by a hidden human operator to serve as a crafting assistant to 20 users, each in an hour-long interaction session that involved multiple handovers of diverse objects used for the crafting task. We then replicate the crafting assistant task to evaluate the autonomous handover model with 20 participants. Further, by comparing the autonomous handover policy and the human operator’s handover strategies, we investigate design implications to improve collaborative handovers. Our main contributions are:
@@
We developed a collaborative handover policy for a mobile manipulator robot, which adapts the handover timing and object transfer location based on a user’s upper body movements using supervised learning;
@@
We compared the objective performance and subjective user perception of autonomous and teleoperated handovers, which informs future HRC research;
@@ II Related Work
As reviewed by Duan et al. [9], recent handover research investigates how different methods for trajectory generation, motion planning, or timing control influence the objective handover performance (e.g., duration or success rate) and subjective human perceptions (e.g., trust or fluency). Existing studies focus on either human-to-robot (H2R) handovers (52.5%) or robot-to-human (R2H) handovers (37.5%), with limited research on bidirectional handovers (10%), i.e., exchanging objects. In H2R handovers, previous work focused on predicting human behaviours or intentions under uncertainty with machine learning models (e.g., [12, 13]). In R2H handovers, previous work focused on enhancing human comfort and increasing the movement efficiency of both parties [14, 15].
@@
Existing handover policies may adopt learning-based, control-based, or analysis-based approaches. For example, Käppler et al. [16] developed an adaptive method for a table-mounted robot arm, which updated the object transfer location based on the location of a user’s hand as observed from an RGB-D camera and proximity sensors on the gripper. This adaptive R2H handover model was compared with a non-adaptive model using a pre-defined object transfer location in a between-subject study, in which the robot handed a cup to a participant in four consecutive repetitions. They found that participants exhibited motor learning and adaptation to the robot’s handover, which necessitated evaluating handover models in repeated episodes. In another example, Kedia et al. [17] developed a transformer-based human intent prediction model pre-trained on human-human collaborative manipulation data and fine-tuned on human-robot data with a teleoperated robot arm. This study incorporated task context in the handover episodes, such as a person and the fixed robot arm taking turns to each pick up one of two objects from a cart to place onto a shared table. However, this work focused on short interaction (3-15s) without direct object transfers between the human and the robot. Zhuang et el. [18] developed an R2H handover model in which a table-mounted robot arm delivered four legs of an IKEA table to a human to assist in assembly. The handover was either triggered by visual-based recognition of human actions, or by voice command (simulated by a supervisor pressing a button after a participant gave the voice command). They evaluated handover time, success rate, and action recognition performance, as well as participants’ perception of the collaboration fluency. This work incorporated collaborative task context in repeated handovers. However, their approach focused on predicting the handover initiation timing, leaving out other timing and location coordination in the whole process.
@@
As shown in the above examples and identified by Duan et al. [9], research on complex long-sequence tasks with handover policies capable of spatial-temporal collaboration is extremely limited. Further, as identified by Ortenzi et al. [1], there is limited work on adaptive handovers incorporating human social and communicative cues beyond hand locations. Moreover, both reviews identified that pre- and post-handover tasks and HRC context were rarely considered in current handover research. Our previous work [11] investigated spatial-temporal adaptation in a mix of H2R, R2H, and bidirectional handovers contextualised in an hour-long HRC task of assembling and painting a wooden birdhouse. While the mobile manipulator robot was teleoperated by a hidden human operator (i.e., Wizard-of-Oz), the FACT HRC dataset collected facilitates the development of a collaborative handover policy by learning from the human operator’s strategies.
III Methodology
@@
We propose to train a collaborative handover policy, using data from a human teleoperator’s adaptive handovers during a crafting task, where multiple purposeful handovers of mixed H2R, R2H, and bidirectional handovers are required [11], as further described in Section III-A. We then evaluate the autonomous handover policy in the same HRC task to understand the task outcomes and users’ subjective experience, as described in Section III-B. We used the Fetch mobile manipulator robot [19] with a 7-DOF arm and a differential drive base, with the robot holding a basket for object delivery and return during handovers. Implementation of the handover model and data collected from our user study (ROS bags, processed csv files, anonymous questionnaire responses) are open-sourced for research purposes. The study protocols were reviewed and approved by Human Research Ethics Committee at Monash University (Project ID 31927).
III-A Developing a collaborative handover model
@@
We trained the autonomous handover policy using the FACT HRC dataset [11], which contains 20 HRC sessions (10 women, 10 men, age 27.05.2, 19 right-handed, 1 left-handed) with 565 handover episodes (57% R2H, 34% H2R, 9% bidirectional) in total. Each HRC session is approximately an hour long with multimodal data collected at time steps of 0.1s. We used the FACT-processed segment of the dataset, which includes non-identifiable csv data of the robot’s status (velocity and coordinates of the base and arm joints, coordinates of the end effector goal for object transfers), the operator’s controls, the facial and upper body keypoints (normalised coordinates of 25 skeletal keypoints with prediction confidences [20]) of participants estimated from an RGB-D camera (OAK-D) positioned next to the participants, and emotion estimations (categorical emotions with intensity, arousal and valence values, inferred from participant’s facial expressions [21]) from this camera and the robot’s onboard camera. Figure 1 provides an overview of the states in each handover episode in the FACT HRC dataset. Within an episode, the human operator was observed to focus their temporal and spatial adaptation, i.e., when and where to perform certain handover actions, at key time points. Thus, instead of continuously predicting the robot’s base and arm actions at every time step, which incurs high computational cost and latency, we segmented the handover episode into four action primitives. In Section IV-B, we discuss our validation of this discretisation of the handover adaptation strategies.
@@
We trained supervised learning policies with the operator’s actions as the ground-truth labels. As illustrated in Figure 1, to achieve temporal adaptation, we trained binary classifiers to predict episode starts (i.e., when the robot starts moving its base towards the participant during action primitive 1), object transfer starts (i.e., when the robot starts stretching out its arm to initiate object transfers during action primitive 2), and handover completions (i.e., when the robot starts tucking its arm and concluding the current handover episode with the participant during action primitive 3). During training, we excluded object transfers with the experimenter (action primitive 4) and focused on learning the operator’s strategies when collaborating with naïve participants. To achieve spatial adaptation, we trained a 3-way classifier to predict the object transfer positions (OTPs) as chosen from the default left (from the robot’s perspective, i.e., to the right of the participant), middle, or right handover goal locations implemented in [11]. Note that the OTP adaptation was performed at both episode starts during action primitive 1 (initial estimation) and object transfer starts during action primitive 2 (updated estimation) as informed by the operator’s strategies observed in [11]. In Section IV, we further discuss our machine learning experiments testing different features and model structures. We implemented the machine learning models using the Python TensorFlow library [22]. The trained classifiers were incorporated with the ROS-based teleoperation framework developed in [11]. Specifically, the classifier outputs determined when and which pre-recorded action primitives the robot would execute, replacing the operator’s controls.
III-B User study evaluation
@@
We replicated the HRC task in [11], where the Fetch mobile manipulator robot served as a crafting assistant by transporting a list of task-relevant objects between a participant who sat at a work desk and an experimenter who sat at a storage desk (Figure 1). The hour-long session was divided into three stages, namely Stage 1 (Preparation, where the participant receives protective equipment, such as a box of tissues and gloves), Stage 2 (Assembly, where the participant receives wooden pieces and glues them together into a birdhouse), and Stage 3 (Painting, where the participant receives brushes and paints to colour the birdhouse to their liking). The list of objects and crafting instructions were displayed on the work and storage desks. The participants were instructed to return objects they no longer need to necessitate a mix of R2H, H2R, and bidirectional handovers.
@@
An observer annotated the handover episodes in terms of quality (Good, Bad, Neutral) and type (H2R, R2H, bidirectional), took observation notes, as well as monitored for program execution and technical issues. ROS bags and processed csv data in the same format as the FACT HRC dataset were collected. The participant filled in an online questionnaire, which collected the demographic information and the participant’s ratings on the robot self-efficacy scale [23] in the pre-study section. After each of the three stages, the participants rated their subjective perception of the robot [24] and the HRC [25]. The robot was paused when the participants filled in the questionnaire. After the whole HRC session, the participants were asked to guess whether the robot was autonomous or teleoperated. They were also interviewed to gather further understanding of their behaviours and experiences.
@@
We conducted the experiments at the Monash Robotics lab. We recruited 20 participants (9 women, 11 men, age 26.34.6, 19 right-handed and 1 left-handed) from the university student and staff population via printed and digital advertisement, direct contact and snowballing. Participants who took part in our previous experiments [11] were excluded from this study. In addition to analysing objective and subjective performance of the handover policy in Section V as a within-subject study, we compared the autonomous handovers with the teleoperated handovers in Section VI as a between-subject study. We hypothesise that the autonomous handover policy will achieve objective and subjective performance similar to human teleoperation.
IV Machine Learning Experiments
@@
IV-A Performance on the FACT HRC dataset
To predict base and arm actions during handovers, we tested features based on the emotion predictions from both cameras and features based on the upper body poses inferred from the OAK-D camera. While combining emotions and poses yielded better performance than poses alone, this improvement was limited due to the low accuracy of the emotion inference. Further, the human operator reported using emotions mainly for evaluating handover quality rather than determining handover strategies [11]. Considering the increased latency of extracting both emotion and pose features, we chose the pose features, i.e., (, , , confidence) of 25 upper body keypoints as inputs of the autonomous handover model.
@@ IV-A Performance on the FACT HRC dataset
In terms of model structure, we incorporated spatial and temporal information in the feature representation and the architecture of the neural networks. We compared Bidirectional Long Short-Term Memory Recurrent Neural Network (BLSTM-RNN) [26], Transformers [27], and Convolutional Neural Network (CNN) [28]. Grid search was used for identifying hyperparameters, as well as the sliding window sizes for padding the input feature vector with preceding steps. We found that the CNN model consistently outperformed BLSTM-RNN and Transformers, which is likely due to spatial relationships of the skeleton keypoints and the relatively small size (65k) of the handover adaptation subset of the FACT HRC dataset. Therefore, we chose the CNN model in our autonomous handover implementation. Our handover classifier includes an input layer, followed by six CNN-1D layers and one dense layer (an architecture inspired by the AlexNet [29], layer sizes [8,8,8,16,16,16,8] from bottom to top), and the output layer. The input feature vectors were padded with a window size of 5 (i.e., including 4 history time steps). The classifier was trained using the Adam optimiser [30] with a learning rate of 0.0001 and a batch size of 8. We adopted early stopping on validation loss with a patience of 20 epochs during training, with a maximum of 100 epochs.
Due to potential individual variances between participants and possible learning effects by the operator and the participants over time, we evaluated the model both with 5-fold CV segmented by participants (i.e., using all episodes from 20% of participants in a testing fold) and 5-fold CV segmented by episodes (i.e., use 20% episodes from each participant in a testing fold). Table I reports average accuracy of the CNN classifiers on the FACT HRC dataset.
As shown in Table I, the timing classifiers (episode start, OTP start, OTP completion) had higher accuracy than the OTP type classifier. The difficulty in OTP type prediction is likely due to two reasons: Firstly, during teleoperation, the human operator was observed to use object layouts to identify a suitable OTP that both aligned with the participant’s behaviours and reduced risks of collision with objects on the desk. As the autonomous policy only used human pose information, the reduced visual input may have limited its OTP prediction. Secondly, the human operator based their timing adaptation strategies mainly with short-term context (e.g., the participant had completed one assembly step and thus was ready for the next piece), while OTP location adaptation strategies involved further understanding of long-term context, such as individual preferences and task progress, which we further discuss in Section VI-C. As the autonomous policy focused on key state transitions within each episode, the lack of global information may have limited its performance.
@@
IV-B Validating action primitives
To verify our model of handover adaptation as fixed action primitives triggered by a set of parameterised state transitions (see Figure 1), we investigated instances when the human operator did not follow these assumptions during teleoperated handovers in the FACT HRC dataset, i.e., when they performed timing adaptation outside of the four specified state transition points and when a non-default OTP was used. We found that the model captured 94.3% of the operator’s adaptations: in 1.2% of episodes, the operator paused during arm reaching; in 2.3% of episodes, the operator paused during arm tucking; in 2.8% of episodes, the operator used non-default OTP, including stopping the end effector early.
V User Study Evaluation Results
We analyse the handover data, questionnaires, observation notes and interviews to evaluate performance of the autonomous handover policy in the HRC crafting task.
V-A Handover episodes
We analysed the processed csv data, including the observer’s in-session annotation of handover type (H2R, R2H, bidirectional) and quality (good, bad, neutral), automatic log of episode count, OTP (left, middle, right of the robot), and handover timing (episode duration, pause at the work space and the storage space). Note that data from P6, P10, P14, P16 were excluded in this analysis due to incomplete manual annotation recording or episode indexing error caused by connection drops. This results in a total of 574 handover episodes from 16 participants.

### v2

@@ Abstract
Robots are increasingly working alongside people, delivering food to patrons in restaurants or helping workers on assembly lines. These scenarios often involve object handovers between the person and the robot. To achieve safe and efficient human-robot collaboration (HRC), it is important to incorporate human context in a robot’s handover strategies. We develop a collaborative handover model trained on human teleoperation data collected in a naturalistic crafting task. To evaluate its performance, we conduct cross-validation experiments on the training dataset as well as a user study in the same HRC crafting task. The handover episodes and user perceptions of the autonomous handover policy were compared with those of the human teleoperated handovers. While the cross-validation experiment and user study indicate that the autonomous policy successfully achieved collaborative handovers, the comparison with human teleoperation revealed avenues for further improvements.
1 Introduction
@@
Object handover is a common task in human-robot collaboration (HRC), in which an item is transferred between a human and a robot [?]. To develop an autonomous human-robot handover policy, existing approaches either react to pre-defined triggers (e.g., specific hand gestures [?]), or train machine learning models using data collected from isolated handover episodes between two people or in human-robot pairs (e.g., [?; ?]). However, using pre-defined triggers imposes learning costs for the person and interrupts the flow of activities, requiring additional physical and mental effort for the person to explicitly command the robot [?]. Further, as handovers are context-dependent [?], training from isolated episodes limits the resulting model’s applicability in handovers that occur in naturalistic contexts with pre- and post-handover tasks.
@@
There has been growing interest in the social-interactive aspects of handovers in addition to the control and trajectory planning aspects, focusing on how a person and a robot may communicate and coordinate during handovers [?; ?]. While these studies investigated communicating information such as intention or hesitation during individual episodes of object transfer, how people and robots collaborate in a complex task involving multiple handovers remains largely under explored. Handovers in a broader task context require providing objects at the right place and the right time, when the person may be engaged in the task and not giving explicit cues to the robot. Further, research on human behaviours in handovers largely focuses on hand locations, while additional informative behaviours such as gaze or whole body movements are under explored [?]. Evaluation of human-robot handover policies is often conducted in a short-term, handover-centred design without incorporating human and task contexts, e.g., focusing on the smoothness of the robot’s motions and trajectories [?]. This results in a gap between a handover policy’s outcomes in isolated episodes compared to when multiple handovers are serving an overall task goal. For example, a policy that emphasises optimising the instantaneous ergonomics risk factors in individual handovers may become over-assistive and reduce worker well-being in repeated interactions [?].
[…]
@@
Therefore, we are motivated to understand human-robot handovers in a naturalistic task context where they serve a collaborative goal. In particular, we develop an autonomous handover policy, as shown in Figure 1, trained with the Functional And Creative Tasks Human-Robot Collaboration dataset (the FACT HRC dataset) [?]. In the FACT HRC dataset, a mobile manipulator robot was teleoperated by a hidden human operator to serve as a crafting assistant to 20 users, each in an hour-long interaction session that involved multiple handovers of diverse objects used for the crafting task. We then replicate the crafting assistant task to evaluate the autonomous handover model with 20 new participants. Further, by comparing the autonomous handover policy and the human operator’s handover strategies, we investigate design implications to improve collaborative handovers. Our main contributions are:
@@
We developed a collaborative handover policy for a mobile manipulator robot, which adapts the handover timing and object transfer location based on a user’s upper body movements;
@@
We compared the objective performance and subjective user perception of autonomous and teleoperated handovers to inform HRC research;
@@ 2 Related Work
As reviewed by Duan et al. [?], research on complex long-sequence tasks with handover policies capable of spatial-temporal collaboration is extremely limited. Recent handover research investigates how different methods for trajectory generation, motion planning, or timing control influence the objective handover performance (e.g., duration or success rate) and subjective human perceptions (e.g., trust or fluency). Existing studies focus on human-to-robot (H2R) or robot-to-human (R2H) handovers, with limited research on bidirectional handovers, i.e., exchanging objects. In H2R handovers, previous work focused on predicting human behaviours or intentions under uncertainty with machine learning models (e.g., [?; ?]). In R2H handovers, previous work focused on enhancing human comfort and increasing the movement efficiency of both parties [?; ?].
@@
Existing handover policies may adopt learning-based, control-based, or analysis-based approaches. For example, Käppler et al. [?] developed an adaptive method for a table-mounted robot arm, which updated the object transfer location based on the location of a user’s hand. This adaptive R2H handover model was compared with a non-adaptive model using a pre-defined object transfer location in four repeated handovers. They found that participants exhibited motor learning and adaptation to the robot’s handover, which necessitated evaluating handover models in repeated episodes. Kedia et al. [?] developed a transformer-based human intent prediction model pre-trained on human-human collaborative manipulation data and fine-tuned on human-robot data with a teleoperated robot arm. They incorporated task context, such as a person and the fixed robot arm taking turns to each pick up one of two objects from a cart to place onto a shared table. However, this work focused on short interaction (3-15s) without direct object transfers between the human and the robot. Zhuang et el. [?] developed an R2H handover model in which a table-mounted robot arm delivered four legs of a table to a human to assist in assembly. They evaluated handover time, success rate, action recognition performance, and participants’ perceptions. This work incorporated collaborative task context in repeated handovers. However, their approach focused on predicting the handover initiation timing, leaving out other timing and location coordination in the whole handover process.
@@
As identified by Ortenzi et al. [?], there is limited work on adaptive handovers incorporating human social and communicative cues beyond hand locations. Moreover, pre- and post-handover tasks and HRC context were rarely considered in current handover research. Our previous work [?] investigated spatial-temporal adaptation in a mix of H2R, R2H, and bidirectional handovers contextualised in an hour-long HRC task of assembling and painting a wooden birdhouse. While the mobile manipulator robot was teleoperated by a hidden human operator (i.e., Wizard-of-Oz), the FACT HRC dataset collected in this work facilitates the development of a collaborative handover policy by learning from the human operator’s strategies.
3 Methodology
@@
To develop a collaborative handover policy capable of spatial and temporal adaptation based on human and task context in purposeful handovers, we used data from a human teleoperator’s adaptive handovers during a crafting task, where multiple purposeful handovers of mixed H2R, R2H, and bidirectional handovers are required [?], as further described in Section 3.1. We then evaluate the autonomous handover policy in the same HRC task to understand the task outcomes and users’ subjective experience, as described in Section 3.2. We used the Fetch mobile manipulator robot [?] with a 7-DOF arm and a differential drive base, with the robot holding a basket for object delivery and return during handovers. Implementation of the handover model and data collected from our user study (ROS bags, processed CSV files, anonymous questionnaire responses) are open-sourced for research purposes. The study protocols were reviewed and approved by Human Research Ethics Committee at Monash University (Project ID 31927).
3.1 Collaborative handover model
@@
We trained the autonomous handover policy using the FACT HRC dataset [?], which contains 20 HRC sessions (10 women, 10 men, age 27.05.2, 19 right-handed, 1 left-handed) with 565 handover episodes (57% R2H, 34% H2R, 9% bidirectional) in total. Each HRC session is approximately an hour long with multimodal data collected at time steps of 0.1s. We used the FACT-processed segment of the dataset, which includes non-identifiable CSV data of the robot’s status (velocity and coordinates of the base and arm joints, coordinates of the end effector goal for object transfers), the operator’s controls, the facial and upper body keypoints (normalised coordinates of 25 skeletal keypoints with prediction confidences [?]) of participants estimated from an RGB-D camera (OAK-D) positioned next to the participants, and emotion estimations (categorical emotions with intensity, arousal and valence values, inferred from participant’s facial expressions [?]) from this camera and the robot’s onboard camera. Figure 1 provides an overview of the states in each handover episode in the FACT HRC dataset. Within an episode, the human operator was observed to focus their temporal and spatial adaptation, i.e., when and where to perform certain handover actions, at key time points. Thus, instead of continuously predicting the robot’s base and arm actions at every time step, which incurs high computational cost and latency, we segmented the handover episode into four action primitives (APs). In Section 4.2, we discuss our validation of this discretisation of the handover adaptation strategies.
@@
We trained supervised learning policies with the operator’s actions as the ground-truth labels. As illustrated in Figure 1, to achieve temporal adaptation, we trained binary classifiers to predict episode starts (i.e., when the robot starts moving its base towards the participant during AP1), object transfer starts (i.e., when the robot starts stretching out its arm to initiate object transfers during AP2), and handover completions (i.e., when the robot starts tucking its arm and concluding the current handover episode with the participant during AP3). During training, we excluded object transfers with the experimenter (AP4) and focused on learning the operator’s strategies when collaborating with naïve participants. To achieve spatial adaptation, we trained a 3-way classifier to predict the object transfer positions (OTPs) as chosen from the default left (from the robot’s perspective, i.e., to the right-hand side of the participant), middle (centre of the work space), or right (left-hand side of the participant) handover goal locations implemented in [?]. The operator chose one of these OTPs to ensure an object can be passed at a location convenient to the participant and suitable for the collaborative task context. Note that the OTP adaptation was performed at both episode starts during AP1 (initial estimation) and object transfer starts during AP (updated estimation) as informed by the operator’s strategies observed in [?]. In Section 4, we further discuss our machine learning experiments testing different features and model structures. We implemented the machine learning models using the Python TensorFlow library [?]. The trained classifiers were incorporated with the ROS-based teleoperation framework developed in [?]. Specifically, the classifier outputs determined when and which pre-recorded APs the robot would execute, replacing the operator’s controls.
3.2 User study evaluation
@@
We replicated the HRC task in [?], where the Fetch mobile manipulator robot served as a crafting assistant by transporting a list of task-relevant objects between a participant who sat at a work desk and an experimenter who sat at a storage desk (Figure 1). The hour-long session was divided into three stages, namely Stage 1 (Preparation, where the participant receives protective equipment, such as a box of tissues and gloves), Stage 2 (Assembly, where the participant receives wooden pieces and glues them together into a birdhouse), and Stage 3 (Painting, where the participant receives brushes and paints to colour the birdhouse to their liking). The list of objects and crafting instructions were displayed on the work and storage desks. The participants were instructed to return objects they no longer need to necessitate a mix of R2H, H2R, and bidirectional handovers.
@@
An observer annotated the handover episodes’ quality (Good, Bad, Neutral) and type (H2R, R2H, bidirectional), took observation notes, and monitored for program execution and technical issues. ROS bags and processed CSV data in the same format as the FACT HRC dataset were collected. The participant filled in a questionnaire, which collected their demographic information and ratings on the robot self-efficacy scale [?] in the pre-study section. After each of the three stages, the participants rated their subjective perception of the robot [?] and the HRC [?]. The robot was paused when the participants filled in the questionnaire. After the whole HRC session, the participants were asked to guess whether the robot was autonomous or teleoperated. They were also interviewed to gather further understanding of their behaviours and experiences.
@@
We conducted the experiments at the Monash Robotics lab. We recruited 20 participants (9 women, 11 men, age 26.34.6, 19 right-handed and 1 left-handed) from the university student and staff population via printed and digital advertisement, direct contact and snowballing. Participants who took part in our previous experiments [?] were excluded from this study. In addition to analysing objective and subjective performance of the handover policy in Section 5 as a within-subject study, we compared the autonomous handovers with the teleoperated handovers in Section 6 as a between-subject study. We hypothesise that the autonomous handover policy will achieve objective and subjective performance similar to human teleoperation.
4 Machine Learning Experiments
@@
4.1 FACT HRC dataset evaluation
To predict base and arm actions during handovers, we tested features based on the emotion predictions from both cameras and features based on the upper body poses inferred from the OAK-D camera. While combining emotions and poses yielded better performance than poses alone, this improvement was limited due to the low accuracy of the emotion inference. Further, the human operator reported using emotions mainly for evaluating handover quality rather than determining handover strategies [?]. Considering the increased latency of extracting both emotion and pose features from two camera feeds, we chose the pose features, i.e., (, , , confidence) of 25 upper body keypoints, as inputs of the autonomous handover model.
@@ 4.1 FACT HRC dataset evaluation
In terms of model structure, we incorporated spatial and temporal information in the feature representation and the architecture of the neural networks. We compared Bidirectional Long Short-Term Memory Recurrent Neural Network (BLSTM-RNN) [?], Transformers [?], and Convolutional Neural Network (CNN) [?]. Grid search was used for identifying hyperparameters, as well as the sliding window sizes for padding the input feature vector with preceding steps. We found that the CNN model consistently outperformed BLSTM-RNN and Transformers, which is likely due to spatial relationships of the skeleton keypoints and the relatively small size (65k) of the handover adaptation subset of the FACT HRC dataset. Therefore, we chose the CNN model in our autonomous handover implementation. Our handover classifier includes an input layer, followed by six CNN-1D layers and one dense layer (an architecture inspired by the AlexNet [?], layer sizes [8,8,8,16,16,16,8] from bottom to top), and the output layer. The input feature vectors were padded with a window size of 5 (i.e., including 4 history time steps). The classifier was trained using the Adam optimiser [?] with a learning rate of 0.0001 and a batch size of 8. We adopted early stopping on validation loss with a patience of 20 epochs during training, with a maximum of 100 epochs.
Due to potential individual variances between participants and possible learning effects by the operator and the participants over time, we evaluated the model both with 5-fold CV segmented by participants (i.e., using all episodes from 20% of participants in a testing fold) and 5-fold CV segmented by episodes (i.e., use 20% episodes from each participant in a testing fold). Table 1 reports average accuracy of the CNN classifiers on the FACT HRC dataset.
As shown in Table 1, the timing classifiers (episode start, OTP start, OTP completion) had higher accuracy than the OTP type classifier. The difficulty in OTP type prediction is likely due to two reasons: Firstly, during teleoperation, the human operator was observed to use object layouts to identify a suitable OTP that both aligned with the participant’s behaviours and reduced risks of collision with objects on the desk. As the autonomous policy only used human pose information, the reduced visual input may have limited its OTP prediction. Secondly, the human operator based their timing adaptation strategies mainly with short-term context (e.g., the participant had completed one assembly step and thus was ready for the next piece), while OTP location adaptation strategies involved further understanding of long-term context, such as individual preferences and task progress, which we further discuss in Section 6.3. As the autonomous policy focused on key state transitions within each episode, the lack of global information may have limited its performance.
@@
4.2 Validating action primitives
To verify our model of handover adaptation as fixed APs triggered by a set of parameterised state transitions (see Figure 1), we investigated instances when the human operator did not follow these assumptions during teleoperated handovers in the FACT HRC dataset, i.e., when they performed timing adaptation outside of the four specified state transition points and when a non-default OTP was used. We found that the model captured 94.3% of the operator’s adaptations: in 1.2% of episodes, the operator paused during arm reaching; in 2.3% of episodes, the operator paused during arm tucking; in 2.8% of episodes, the operator used non-default OTP, including stopping the end effector early.
5 User Study Evaluation Results
We analyse the handover data, questionnaires, observation notes and interviews to evaluate the autonomous handover policy in the HRC crafting task.
5.1 Handover episodes
We analysed the processed CSV data, including the observer’s in-session annotation of handover type (H2R, R2H, bidirectional) and quality (good, bad, neutral), automatic log of episode count, OTP (left, middle, right of the robot), and handover timing (episode duration, pause at the work space and the storage space). Note that data from P6, P10, P14, P16 were excluded here due to incomplete annotation or episode indexing error caused by connection drops. This results in a total of 574 handover episodes from 16 participants.